OpenStudy (anonymous):

since Iq=Uq-Lq=8.2, median is given 150lb , how we can find lower quartile?

OpenStudy (campbell_st):

well it is a value between 150 and 141.8 this is because you have no way of knowing the data points... as an example..if the data set was 150, 150, 150, 158.2 170 median is 150 and the lower quartile is 150...

OpenStudy (anonymous):

@campbell_st Thanx

OpenStudy (anonymous):

@campbell_st how to find 95th percentile of this question

OpenStudy (campbell_st):

well divide the data into 10 equal parts... the 95 percentile is half way between the 90th and 100% percentile..

OpenStudy (anonymous):

@campbell_st One of the petals is 0.2 cm wider than another. Regression says that the wider petal is estimated to be ____________ cm longer than the other

OpenStudy (campbell_st):

sorry... its an age since I did regression analysis... you would need the regression equation and substitute x + 0.2 to estimate the length of the leaf... I'd expect

OpenStudy (anonymous):

Q3-Q1= SD(z3-z1) 8.2 = SD(0.67-(-0.67)) SD=8.2/1.34 SD=6.12